OpenStudy (anonymous):

Find i (the rate per period) and n (the number of periods) for the following annuity. Quarterly deposits of $1,500 are made for 3 years into an annuity that pays 7.5% compounded quarterly. i=? n=? I got 0.075 and 12, but am I right?

OpenStudy (amistre64):

i is off

OpenStudy (amistre64):

7.5 is a yearly interest amount, only a quarter of it is applied in one quarter

OpenStudy (amistre64):

4 times a year for 3 years is fine for n

OpenStudy (anonymous):

O.k. ty

OpenStudy (amistre64):

i might have thought to hard on i you are trying to fill in a formula and if it has i/4 then you are fine with that too

OpenStudy (amistre64):

\[P(1+\frac in)^{nt}\] if this is your formula, n=4, and i = .075

OpenStudy (amistre64):

if your formula is:\[P(1+i)^n\] then i = .075/4 and n=12 so its best to determine what exactly it is asking for.

OpenStudy (anonymous):

After rereading it, I concur, I think P(1+i)^n is correct.

OpenStudy (anonymous):

\[3000=500*1.02^n-1/.02\] solve for n

OpenStudy (amistre64):

i cant verify that, so i defer to your judgement on it :)

OpenStudy (amistre64):

im not sure i can read your equation correctly ... what is the last term?

OpenStudy (anonymous):

\[FV=PMT(1+i)^n-1 \over i\]

OpenStudy (anonymous):

where FV= 3000 PMT = 500 i=.02

OpenStudy (anonymous):

Its just that I keep getting different answers when I solve for n

OpenStudy (amistre64):

\[F=\frac{P(1+i)^n-1}{i}\] \[Fi+1=P(1+i)^n\] \[\frac{Fi+1}{P}=(1+i)^n\] log it to solve for n

OpenStudy (anonymous):

ok thanks

OpenStudy (amistre64):

yep